A Bibliography of Publications of Diomidis Spinellis

Diomidis Spinellis University of the Aegean

GR-83200 Karlovasi Greece

Tel: ?n/a? FAX: ?n/a?

E-mail: dspin@aegean.gr (Internet)

01 March 2018 Version 1.15

Abstract

This bibliography records publications of Diomidis Spinellis.

Title word cross-reference

$39.99 [82].

0 [82]. 0-321-28608-1 [82].

1 [82]. 1947 [27].

5th [120].

’94 [120].

Abstraction [84]. Abusing [99]. Access [41, 25]. ACM [76]. Adaptive [43]. Addison [82]. Addison-Wesley [82]. Addressing [28]. Advancement [80]. Advisories [67]. Affecting [64]. Age [52]. Agile [82, 125]. Agility [115]. Allocation [33, 48]. Already [76]. Aluminum [23]. Analysis [29, 101]. Annealing [34, 33, 48]. Annual [123, 124]. Antikythera [95]. api [87, 63, 7]. Appliance [37]. Applications [49, 3, 6, 7, 8, 9, 10, 11, 12, 20, 21, 120, 65]. Approach [34, 67, 43, 15, 33, 109, 48]. Architectural [13]. Architecture [52, 15, 45, 102]. architectures [119]. Art [72]. Asset [81]. Athens [120]. Attract [76]. Automatic [38]. Automation [23].

Barbara [122]. Based [41, 46, 109]. Basic [103]. Beautiful [102]. Bias [76]. Book [20, 21]. Bookshelf [62, 52, 43, 44, 68, 85]. Boston [124]. bounded [59]. bounded-length [59]. Brain [66]. browser [111]. Bruce [20, 21]. Buffer [33, 48]. Buffering [41]. Bug [77]. builder [54]. Builders [98]. Building [5]. Bullets [90]. Busters [77].

C [4, 111]. CA [123]. California [121, 122]. Call [18]. Cascade [30, 34]. CD [22]. CD-ROM [22]. Change} [43]. channel [63]. Character [26, 27]. Checking [4].

Choosing [31, 78, 116]. Classes [15, 16]. Clothes [36]. Code [60, 110, 35]. Coding [50]. Collaborative [43, 96]. commercial [63]. Communication [103]. complete [59]. Complex [43]. Component [39]. Components [116, 54]. Composition [13]. Compression [9]. Computer [36, 95]. Computing [66]. Conference [121, 122, 123, 124, 119, 120]. content [61]. Control [10, 73, 74, 43]. Controlling [12]. Controls [37]. Conversation [31]. Copley [124]. Copying [17]. Copyright [76]. Correction [34]. Correspondence [42]. Course [76]. Court [22]. Cover [76]. Covering [76]. Cracking [86]. Cryptographic [29]. CS [66, 76]. CScout [111]. Curing [53].

Database [120, 14]. Dear [69]. Debuggers [79]. decay [55]. Decisions [22]. Declarations [4]. Declarative [40, 42, 56]. Design [13, 14, 20, 21, 101, 51]. Detection [34]. Development [43, 8, 64, 82, 63, 125]. DEXA [120]. Diego [123]. Difficult [106]. DIKAIO [22]. dimensional [13]. Disband [66]. Discipline [52]. discrete [101]. Disks [112]. Distributed [29, 30]. distribution [61]. Documentation [110]. Domain [108, 122, 51]. Domain-Specific [108, 122, 51]. Don’t [76]. Double [41]. Double-Buffering [41]. Downturns [66]. Dr [20, 21]. Drawing [105]. Drivers [115]. DTV [63].

Editor [69]. Editors [64, 108]. edoC [117]. Education [76]. Electronic [26, 27]. elytS [117]. Embedded [52]. Embrace [43]. Encoding [26, 27]. Engineering [32]. environment [54]. Equal [83]. Etiquette [103]. Everywhere [114]. Excogitate [39]. Exercise [5]. expert [120]. Explained [43]. Exploit [39]. Explore [39, 45]. expression [93]. extending [63]. eXtreme [43]. failures [55]. Faking [118]. Farewell [112]. Fear [50]. Field [53]. File [9, 58]. FIRE [93]. FIRE/J [93]. Flaws [29]. FOCUS [108]. Forgotten [72]. Forum [66, 76]. framework [87]. Frameworks [79]. Francisco [121]. FREENIX [124]. Function [66]. Functions [2]. Fundamentals [76]. Future [76].

Generation [38]. generative [93]. Gentle [52]. Government [66]. Greece [120]. Greek [27, 26]. Guest [64, 108]. GUI [54]. GUI-builder [54].

Hacker [66]. Hammer [108]. Hand [71]. Handling [67, 11]. Haskell [1, 5]. Here [76]. High [22, 20, 21]. High-Speed [20, 21]. Hot [23]. Hotel [124].

identification [59]. Images [11]. Implementation [3, 5, 14, 24, 31, 1]. Implementing [5, 6, 7, 8, 9, 10, 11, 12]. Increase [76]. Integrated [23]. Integrity [47]. interactive [63]. Interface [10, 63]. international [119, 120]. Internet [22, 52]. Introduction [52, 6, 44, 64, 108]. Irrelevant [70]. ISBN [82]. Issues [28].

J [93]. January [121]. Java [70]. Jetpack [92]. Job [104]. Journeyman [44]. June [123, 124].

Kevin [82]. Kinds [108]. knowledge [96]. Krell [20, 21].

Language [5, 15, 24, 78, 1]. Languages [32, 70, 122, 35, 119, 51]. laws [94]. Legal [14]. length [59]. level [107]. Lightweight [32]. line [45]. Lines [33, 101, 48]. Link [4]. Linkage [2]. Logging [18, 79]. Loyal [36].

Machine [12, 25]. Mail [26, 27]. Makes [70]. Management [43]. Managing [43, 109]. Mandate [76]. March [119]. market [109]. market-based [109]. Marriott [124]. Massachusetts [124]. Master [44]. Matching [42, 40]. Maturing [52]. Mechanism [16, 47, 95]. Media [10]. Meets [46]. Members [76]. Memory [83]. Messages [26, 27]. Metaprogramming [97]. Methods [9, 20, 21]. Milling [23]. Mining [39]. Misteach [76]. Misuse [76]. mobile [35]. Modeling [108]. Models [56]. Modest [53]. Modular [45]. Most [106]. Moving [11]. MS [25]. MS-Access [25]. Multimedia [3, 6, 7, 8, 9, 10, 11, 12]. Multiparadigm [15, 16, 24]. Multitasking [20, 21]. My [92]. Mysteries [90].

Nails [108]. Need [108]. networks [65]. No [76]. Notable [51]. NP [59]. NP-complete [59].

Object [15, 16, 24, 39]. Object-Oriented [24]. October [122]. Open [29, 30, 80, 116, 64]. Operating [18, 107]. Opposition [36]. Optimal [41]. Optimization [42, 40, 45]. optimizing [93]. organization [96]. Organized [58]. Oriented [24]. Other [90]. Outwit [46].

Palmtop [37]. Paper [88]. Paradigms [15, 16]. Parallels [66]. Part [106, 101, 73]. Pattern [42, 40]. patterns [51]. Peephole [42, 40]. peer [61, 109, 65]. peer-to-peer [61, 109, 65]. Peripheral [41]. Perl [31]. Perspective [82, 95, 125]. Phobia [53]. Pipe [41]. Pipe-Based [41]. Place [124]. platform [63]. Portability [81]. Posse [66]. Power [94]. Practical [52]. Pragmatic [44]. Problem [30, 34]. problems [35]. Proceedings [121, 122, 124, 119, 120]. Process [43, 44]. Production [76, 33, 101, 45, 48]. Professional [80, 82]. Program [13, 100]. Programmable [37]. Programmer [44]. Programming [35, 119, 43, 7, 15, 16, 46, 78, 93, 54]. Project [81]. Proposal [53]. Protecting [17]. Protocol [29]. Protocols [29]. Pruning [58]. Public [53].

Quality [43].

Rational [97]. Reading [60]. Reduce [50]. refactoring [111]. references [55]. Reflection [47]. Reflections [57]. regular [93]. Related [29]. Reliable [33, 59, 48]. Reporting [67]. Response [66]. Reuse [86]. Review [30, 20, 21, 82]. revisited [57]. RFC [27]. risk [109]. risks [57]. ROM [22]. Room [76].

safe [2]. San [121, 123]. Sandboxing [49]. Santa [122]. Science [95]. Scope [76]. Scorecard [67]. Screen [25]. Scripting [70]. searches [93]. Security [28, 67, 104, 65]. September [120]. Sets [58]. Shell [31]. Silver [90]. Simplicity [76]. Simulated [34, 33, 48]. Small [38]. Software [52, 43, 17, 32, 44, 47, 64, 82, 86, 98, 113, 94, 125]. Some [83]. Source [80, 116, 64]. Specific [108, 122, 51]. Specification [56]. Speed [20, 21]. Spy [89]. Start [106]. static [87]. Storage [9]. Stress [76]. String [42, 40]. Structuring [16]. Subjects [13]. survey [61]. Sustainable [125, 82]. Switzerland [119]. System [76, 18, 119, 107]. Systems [29, 30, 52, 74, 120, 43].

Tate [82]. Taxonomy [29]. Team [44]. Tech [66]. Technical [42, 103, 123, 124]. technologies [61]. Technology [28, 24, 39]. TEKTON [13]. television [63]. Text [14, 38]. Threats [28]. Three [13]. Three-dimensional [13]. Time [4]. Tool [5, 18, 46, 72]. Tool-Based [46]. Toolkit [25]. Tools [29, 8, 19, 32, 38, 71, 75, 91, 105, 106, 54]. Trace [18]. Track [124]. Tracks [113]. Trade [105, 106]. Transactional [52]. Transactions [18, 109, 107]. trust [57]. trusting [57]. Type [2]. Type-safe [2]. Types [83].

UML [76, 114]. Unauthorised [17]. Unified [15]. UNIX [19, 46, 54, 75]. USA [121, 123, 124]. Use [91]. User [63, 107]. User-level [107]. Using [41, 42, 99, 116, 40].

Variables [2]. Variation [84]. Vehicle [31]. Verification [47, 87]. Version [73, 74]. Video [12]. View [13]. viewpoint [35]. virtual [63]. viruses [59]. visual [54]. Vulnerability [30, 34].

Way [100]. Web [28, 55]. Wesley [82]. Where [92]. Wide [28]. Windows [20, 21, 7, 19, 46]. Winter [121]. Working [75]. World [46, 28]. Writing [60, 72]. Wux [19].

XML [99].

Zurich [119].

References

Spinellis:1990:IHL

[1] Diomidis Spinellis. An implementation of the Haskell language. Master’s thesis, Imperial College, London, UK, June 1990.

Spinellis:1991:TSL

[2] Diomidis Spinellis. Type-safe linkage for variables and functions. ACM SIG-PLAN Notices, 26(8):74–79, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://kerkis.

math.aegean.gr/~dspin/pubs/jrnl/
1991-SIGPLAN-CType/html/tsl.html.

Spinellis:1992:IMA

[3] Diomidis Spinellis. The implementation of multimedia applications. In Multimedia: Theory and Experience, pages 23–37. Greek Computer Society, Athens, Greece, December

1992.  URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1992-
EPY-Multimedia-SM/html/epy-mlm2.

html. In Greek.

Spinellis:1993:CCD

[4] Diomidis D. Spinellis. Checking C declarations at link time. The Journal of C Language Translation, 4(3): 238–249, March 1993. ISSN 1042-

5721.  URL http://kerkis.math.
aegean.gr/~dspin/pubs/jrnl/1993-
JCLT-CType/html/tsl.html.

Spinellis:1993:IHL

[5] Diomidis Spinellis. Implementing Haskell: Language implementation as a tool building exercise. Structured Programming, 14(1):37–48, 1993. CO-DEN STPGEM. ISSN 0935-1183.

URL http://kerkis.math.aegean.
gr/~dspin/pubs/jrnl/1993-StrProg-
Haskell/html/exp.html.

Spinellis:1993:IMAa

[6] Diomidis Spinellis. Implementing multimedia applications: Introduction. Win magazine, pages 86–89, March

1993.  URL http://kerkis.math.
aegean.gr/~dspin/pubs/trade/1993-
Winmag/fastwin1/html/win1.html. In

Greek.

Spinellis:1993:IMAb

[7] Diomidis Spinellis. Implementing multimedia applications: Programming the Windows API. Win magazine, pages 64–68, April 1993.

URL http://kerkis.math.aegean.
gr/~dspin/pubs/trade/1993-Winmag/
fastwin2/html/win2.html. In Greek.

Spinellis:1993:IMAc

[8] Diomidis Spinellis. Implementing multimedia applications: Development tools. Win magazine, pages 65–68,

June 1993. URL http://kerkis.math.
aegean.gr/~dspin/pubs/trade/1993-
Winmag/fastwin3/html/win3.html. In

Greek.

Spinellis:1993:IMAd

[9] Diomidis Spinellis. Implementing multimedia applications: File storage and compression methods. Win magazine, pages 80–83, July 1993.

URL http://kerkis.math.aegean.
gr/~dspin/pubs/trade/1993-Winmag/
fastwin4/html/win4.html. In Greek.

Spinellis:1993:IMAe

[10] Diomidis Spinellis. Implementing multimedia applications: The media control interface. Win magazine, pages 89–92, September 1993.

URL http://kerkis.math.aegean.
gr/~dspin/pubs/trade/1993-Winmag/
fastwin5/html/win5.html. In Greek.

Spinellis:1993:IMAf

[11] Diomidis Spinellis. Implementing multimedia applications: Handling moving images. Win magazine, Octo-

ber 1993. URL http://kerkis.math.
aegean.gr/~dspin/pubs/trade/1993-
Winmag/fastwin6/html/win6.html. In

Greek.

Spinellis:1993:IMAg

[12] Diomidis Spinellis. Implementing multimedia applications: Controlling the video machine. Win magazine, Novem-

ber 1993. URL http://kerkis.math.
aegean.gr/~dspin/pubs/trade/1993-
Winmag/fastwin7/html/win7.html. In

Greek.

Spinellis:1993:TPC

[13] Diomidis Spinellis. TEKTON: A program for the composition, design, and three-dimensional view of architectural subjects. In 4th Panhellenic Informatics Conference, volume I, pages 361–372. Greek Computer Society, Patras, Greece, Decem-

ber 1993. URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1993-
EPY-Tekton/html/tekton.html.  In

Greek.

Spinellis:1994:DIL

[14] Diomidis Spinellis. The design and implementation of a legal text database. In Karagiannis [120], pages 339–348. CODEN LNCSD9. ISBN 0-387-58435-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3

D35843 1994. URL http://kerkis.
math.aegean.gr/~dspin/pubs/conf/
1994-DEXA-Dikaio/html/dexa.html.

Spinellis:1994:LAP

[15] Diomidis Spinellis, Sophia Drossopoulou, and Susan Eisenbach. Language and architecture paradigms as object classes: A unified approach towards multiparadigm programming. In Gutknecht [119], pages 191–207. CO-DEN LNCSD9. ISBN 0-387-57840-4 (New York), 3-540-57840-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P79 1994.

DM72.00. URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1994-
PLSA-Multipar/html/paper.html.

Spinellis:1994:PPO

[16] Diomidis Spinellis. Programming Paradigms as Object Classes: A Structuring Mechanism for Multiparadigm Programming. PhD thesis, Imperial College of Science, Technology and Medicine, London, UK, February 1994.

Spinellis:1994:PSU

[17] Diomidis Spinellis. Protecting software from unauthorised copying. Win magazine, pages 75–77, March 1994.

URL http://kerkis.math.aegean.
gr/~dspin/pubs/trade/1993-Winmag/
winlock/html/winlock.html.  In

Greek.

Spinellis:1994:TTL

[18] Diomidis Spinellis. Trace: A tool for logging operating system call transactions. Operating Systems Review, 28(4):56–63, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (elec-

tronic). URL http://kerkis.math.
aegean.gr/~dspin/pubs/jrnl/1994-
SIGOS-Trace/html/article.html.

Spinellis:1994:WUT

[19] Diomidis D. Spinellis. Wux: UNIX tools under Windows. In USENIX Association [121], pages 325–336. ISBN 1-880446-58-8. LCCN QA 76.76 O63

U84 1994. URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1994-
Usenix-Wux/html/wux.html;   http:
//www.usenix.org/publications/library/
proceedings/sf94/spin.html.

Spinellis:1995:BRHa

[20] Diomidis Spinellis. Book review: High-Speed Windows Applications: Multitasking Design Methods by Dr. Bruce E. Krell. ACM SIGPLAN Notices, 30(4):11, April 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1995:BRHb

[21] Diomidis Spinellis. Book review: High-Speed Windows Applications: Multitasking Design Methods, by Dr. Bruce E. Krell. ACM SIGPLAN Notices, 30(7):4, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1995:DHC

[22] Diomidis Spinellis. DIKAIO: High court decisions on the internet and on CD-ROM. In First Greek Technical Chamber Informatics Conference: The Information Society, pages 546–555. Technical Chamber of Greece, Athens, Greece, December 1995. URL http:

//kerkis.math.aegean.gr/~dspin/
pubs/conf/1995-TEE-InfSoc-Dikaio/
html/dikaio.html. In Greek.

Spinellis:1995:IAH

[23] Diomidis Spinellis. Integrated aluminum hot milling automation. In Modern Automated Control Technologies, pages 218–221. Technical Chamber of Greece, Athens, Greece, Decem-

ber 1995. URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1995-
TEE-Control-OIS/html/teeconf1.html.

In Greek.

Spinellis:1995:OOT

[24] Diomidis Spinellis, Sophia Drossopoulou, and Susan Eisenbach. Objectoriented technology in multiparadigm language implementation. Journal of Object Oriented Programming, 8(1):33–38, March/April 1995.

CODEN JOOPEC. ISSN 0896-

8438.  URL http://kerkis.math.
aegean.gr/~dspin/pubs/jrnl/1995-
JOOP-Multipar/html/paper.html.

Spinellis:1995:SMM

[25] Diomidis Spinellis. The Screen Machine MS-Access toolkit. Win magazine, Jan-

uary 1995. URL http://kerkis.math.
aegean.gr/~dspin/pubs/trade/1993-
Winmag/smdb/html/smdb.html.  In

Greek.

Spinellis:1996:GCE

[26] Diomidis Spinellis. Greek character encoding for electronic mail messages. Network Information Center, Request for Comments 1947, May 1996.

URL http://kerkis.math.aegean.
gr/~dspin/pubs/TR/RFC-1947/html/
RFC1947.html. RFC-1947.

Spinellis:1996:RGC

[27] D. Spinellis. RFC 1947: Greek character encoding for electronic mail messages, May 1996. URL ftp://

ftp.internic.net/rfc/rfc1947.txt;
ftp://ftp.math.utah.edu/pub/rfc/

rfc1947.txt. Status: INFORMA-TIONAL.

Gritzalis:1997:ATS

[28] Stefanos Gritzalis and Diomidis Spinellis. Addressing threats and security issues in World Wide Web technology. In Proceedings CMS ’97 3rd IFIP TC6/TC11 International joint working Conference on Communications and Multimedia Security, pages 33–46. IFIP, Chapman and Hall, Ltd., London, UK, September 1997. URL http://kerkis.

math.aegean.gr/~dspin/pubs/conf/
1997-CMS-WebSec/html/w3sec.html.

Gritzalis:1997:CPO

[29] Stefanos Gritzalis and Diomidis Spinellis. Cryptographic protocols over open distributed systems: A taxonomy of flaws and related protocol analysis tools. In 16th International Conference on Computer Safety, Reliability and Security: SAFECOMP ’97, pages 123–137. European Workshop on Industrial Computer Systems: TC-7, Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., Septem-

ber 1997. URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1997-
SafeComp-Formal/html/doc.html.

Gritzalis:1997:CVP

[30] Stefanos Gritzalis and Diomidis Spinellis. The cascade vulnerability problem for open distributed systems: A review. In IFIP TC5 WG5.4 3rd International Conference on Reliability, Quality & Safety of Software-Intensive Systems, ENCRESS ’97, pages 175–185. IFIP, Chapman and Hall, Ltd., London, UK,

May 1997. URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1997-
Encress-Cascade/html/casc.html.

Spinellis:1997:CAP

[31] Diomidis Spinellis and Rob Kolstad. A conversation about Perl and the shell: Choosing the implementation vehicle. ;login: the USENIX Association newsletter, 22(5):25–31, June 1997. CODEN LOGNEM. ISSN

1044-6397. URL http://kerkis.math.
aegean.gr/~dspin/pubs/trade/1997-
login-Perl/html/shperl.html.

Spinellis:1997:LLS

[32] Diomidis Spinellis and V. Guruprasad. Lightweight languages as software engineering tools. In USENIX [122], pages 67–76. ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http:

//kerkis.math.aegean.gr/~dspin/
pubs/conf/1997-DSL-Lightweight/
html/paper.html;   http://www.usenix.
org/publications/library/proceedings/
dsl97/spinellis.html.

Spinellis:1997:SAA

[33] Diomidis Spinellis and Hrisoleon T. Papadopoulos. A simulated annealing approach for buffer allocation in reliable production lines. In International Workshop on Performance Evaluation and Optimization of Production Lines, pages 365–375. University of the Aegean, Department of Mathematics, Samos, Greece, May

1997.  URL http://kerkis.math.
aegean.gr/~dspin/pubs/conf/1997-
Perf-Anneal/html/anneal.html.

Gritzalis:1998:CVP

[34] Stefanos Gritzalis and Diomidis Spinellis. The cascade vulnerability problem: The detection problem and a simulated annealing approach for its correction. Microprocessors and Microsystems, 21(10):621–628, April 1998. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://kerkis.math.
aegean.gr/~dspin/pubs/jrnl/1997-
uProcessors-Cascade/html/casc.html.

Gritzalis:1998:PLM

[35] Stefanos Gritzalis, George Aggelis, and Diomidis Spinellis. Programming languages for mobile code: A problems viewpoint. In Proceedings of the First International Network Conference INC ’98. IEE, Internet Research, Plymouth, UK, July 1998. URL http://kerkis.

                                        math.aegean.gr/~dspin/pubs/conf/
                                        1998-INC-Mobcode/html/inc98.htm.

Spinellis:1998:LOC

[36] Diomidis Spinellis. Loyal opposition: The computer’s new clothes. IEEE Software, 15(6):14–??, November/December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

                                        URL http://dlib.computer.org/so/
                                        books/so1998/pdf/s6014.pdf.

Spinellis:1998:PPA

[37] Diomidis Spinellis. Palmtop programmable appliance controls. Personal Technologies, 2(1), March 1998.

                                        URL http://kerkis.math.aegean.
                                        gr/~dspin/pubs/jrnl/1998-PeTe-Remote/
                                        html/article.html.

Spinellis:1998:STA

[38] Diomidis Spinellis. Small tools for automatic text generation. ;login: the USENIX Association newsletter, 23(5):44–47, August 1998. CODEN LOGNEM. ISSN 1044-6397. URL http:

                                        //kerkis.math.aegean.gr/~dspin/
                                        pubs/trade/1998-Login-TextTools/
                                        html/paper.html;   http://www.usenix.
                                        org/publications/login/1998-8/tools.
                                        html.

Spinellis:1999:COT

[39] Diomidis Spinellis. Component and object technology: Explore, excogitate, exploit: Component mining. Computer, 32(9):114–116, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/co/
                                        books/co1999/pdf/r9114.pdf.

Spinellis:1999:DPO

[40] Diomidis Spinellis. Declarative peephole optimization using string pattern matching. ACM SIGPLAN Notices, 34(2): 47–50, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1999:OPA

[41] Diomidis Spinellis. Optimal peripheral access using pipe-based doublebuffering. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.usenix.org/
publications/login/1999-8/features/
peripheral.html.

Spinellis:1999:TCD

[42] Diomidis Spinellis. Technical correspondence: Declarative peephole optimization using string pattern matching. ACM SIGPLAN Notices, 34(2): 47–51, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
s/Spinellis:Diomidis.html.

Moitra:2000:BAS

[43] Deependra Moitra, Diomidis Spinellis, and Robert Bruce Kelsey. Bookshelf: Adaptive Software Development: A Collaborative Approach to Managing Complex Systems; eXtreme Programming Explained: Embrace Change; Software Process Quality: Management and Control. IEEE Software, 17(4): 112–115, July/August 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://dlib.

                                        computer.org/so/books/so2000/pdf/
                                        s4112.pdf.

Spinellis:2000:BPP

[44] Diomidis Spinellis and Gerry Coleman. Bookshelf: The Pragmatic Programmer: From Journeyman to Master; Introduction to the Team Software Process. IEEE Software, 17(6):108–110, November/December 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://dlib.

                                        computer.org/so/books/so2000/pdf/
                                        s6108.pdf.

Spinellis:2000:MPL

[45] Diomidis D. Spinellis and Chrissoleon T. Papadopoulos. Modular production line optimization: the exPLOre architecture. Math. Probl. Eng., 6(6):527–541, 2000. ISSN 1024-123X.

Spinellis:2000:OUT

[46] Diomidis D. Spinellis. Outwit: UNIX tool-based programming meets the Windows world. In USENIX [123], pages 149–158. ISBN 1-880446-22-7. LCCN ???? URL http://

                                        www.usenix.org/events/usenix2000/
                                        general/spinellis.html.

Spinellis:2000:RMS

[47] Diomidis Spinellis. Reflection as a mechanism for software integrity verification. ACM Transactions on Information and System Security, 3(1):51–62, February 2000. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Spinellis:2000:SAA

[48] Diomidis D. Spinellis and Chrissoleon T. Papadopoulos. A simulated annealing approach for buffer allocation in reliable production lines. Ann. Oper. Res., 93:373–384, 2000. ISSN 0254-5330 (print), 1572-9338 (electronic). Performance evaluation and optimization of production lines (Samos Island, 1997).

Prevelakis:2001:SA

[49] Vassilis Prevelakis and Diomidis Spinellis. Sandboxing applications. In USENIX [124], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001.  URL http://www.
usenix.org/publications/library/
proceedings/usenix01/freenix01/
prevelakis.html.

Spinellis:2001:FCH

[50] Diomidis Spinellis. Fear of coding, and how to reduce it. Computer, 34(8): 100, 98–99, August 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r8100.pdf;   http://www.computer.
org/computer/co2001/r8100abs.htm.

Spinellis:2001:NDP

[51] Diomidis Spinellis. Notable design patterns for domain-specific languages. The Journal of Systems and Software, 56 (1):91–99, February 1, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/29/11/
57/24/32/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/29/11/
57/24/32/article.pdf.

Larrabee:2002:BSA

[52] Robert C. Larrabee, Diomidis Spinellis, and Angelo Bellotti. Bookshelf: Software architecture: A maturing discipline; A gentle introduction to embedded systems; practical transactional systems in the Internet Age. IEEE Software, 19(1):100–102, January/February 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

                                        tronic). URL http://computer.org/
                                        software/so2002/s1100abs.htm;
                                        http://dlib.computer.org/so/books/
                                        so2002/pdf/s1100.pdf.

Spinellis:2002:MPC

[53] Diomidis Spinellis. A modest proposal for curing the public field phobia. ACM SIGPLAN Notices, 37(4):54–56, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:2002:UTV

[54] Diomidis Spinellis. Unix tools as visual programming components in a GUI-builder environment. Software—Practice and Experience, 32 (1):57–71, January 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/88013123/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=88013123&PLACEBO=
                                        IE.pdf.

Spinellis:2003:DFW

[55] Diomidis Spinellis. The decay and failures of Web references. Communications of the ACM, 46(1):71–77, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spinellis:2003:DSM

[56] Diomidis Spinellis. On the declarative specification of models. IEEE Software, 20(2):96, 94–95, March/April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.org/so/
books/so2003/pdf/s2096.pdf.

Spinellis:2003:IRR

[57] Diomidis Spinellis. Inside risks: Reflections on trusting trust revisited. Communications of the ACM, 46(6): 112, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spinellis:2003:OPF

[58] Diomidis D. Spinellis. Organized pruning of file sets. ;login: the USENIX Association newsletter, 28(3):??, June 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.usenix.org/
publications/login/2003-06/pdfs/
spinellis.pdf.

Spinellis:2003:RIB

[59] Diomidis Spinellis. Reliable identification of bounded-length viruses is NPcomplete. IEEE Transactions on Information Theory, 49(1):280–284, 2003. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Spinellis:2003:RWC

[60] Diomidis Spinellis. Reading, writing, and code. ACM Queue: Tomorrow’s Computing Today, 1(7):84–89, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Androutsellis-Theotokis:2004:SPP

[61] Stephanos Androutsellis-Theotokis and Diomidis Spinellis. A survey of peerto-peer content distribution technologies. ACM Computing Surveys, 36 (4):335–371, December 2004. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Berzal:2004:B

[62] Fernando Berzal and Diomidis Spinellis. Bookshelf. IEEE Software, 21(3):104– 105, May/June 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chorianopoulos:2004:UID

[63] Konstantinos Chorianopoulos and Diomidis Spinellis. User interface development for interactive television: extending a commercial DTV platform to the virtual channel API. Computers and Graphics, 28(2):157–166, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Spinellis:2004:GEI

[64] Diomidis Spinellis and Clemens Szyperski. Guest Editors’ introduction: How is Open Source affecting software development? IEEE Software, 21(1): 28–33, January/February 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

                                        //csdl.computer.org/comp/mags/so/
                                        2004/01/s1028.pdf.

Vlachos:2004:SAP

[65] Vasileios Vlachos, Stephanos Androutsellis-Theotokis, and Diomidis Spinellis. Security applications of peer-to-peer networks. Computer Networks (Amsterdam, Netherlands: 1999), 45(2):195– 205, June 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Crawford:2005:FWC

[66] Diane Crawford, Diomidis Spinellis, Peter J. Denning, Henry Baker, Peter J. Denning, Alex Simonelis, David Patterson, and Jeff Johnson. Forum: Why computing parallels brain function; the government’s response to tech and CS downturns; disband the hacker posse. Communications of the ACM, 48(10):11–13, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lekkas:2005:HRS

[67] Dimitrios Lekkas and Diomidis Spinellis. Handling and reporting security advisories: A scorecard approach. IEEE Security & Privacy, 3(4):32–41, July/ August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Spinellis:2005:B

[68] Diomidis Spinellis, John R. Dance, David Arthur Eatough, Kevin C. Desouza, and Yukika Awazu. Bookshelf. IEEE Software, 22(1):94–97, January/February 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//csdl.computer.org/comp/mags/so/
2005/01/s1094.pdf.

Spinellis:2005:DE

[69] Diomidis Spinellis. Dear editor. IEEE Software, 22(2):14–15, March/April 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:JMS

[70] Diomidis Spinellis. Java makes scripting languages irrelevant? IEEE Software, 22(3):70–71, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:TH

[71] Diomidis Spinellis. The tools at hand. IEEE Software, 22(1):10–12, January/ February 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:TWF

[72] Diomidis Spinellis. Tool writing: A forgotten art? IEEE Software, 22(4):9–11, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:VCP

[73] Diomidis Spinellis. Version control, Part I. IEEE Software, 22(5):107, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

                                        //csdl.computer.org/comp/mags/so/
                                        2005/05/s5107.pdf.

Spinellis:2005:VCS

[74] Diomidis Spinellis. Version control systems. IEEE Software, 22(5):108– 109, September/October 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2005:WUT

[75] Diomidis Spinellis. Working with Unix tools. IEEE Software, 22(6):9–11, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Crawford:2006:FSF

[76] Diane Crawford, Umesk Panchaksharaiah, L. Peter Deutsch, Kallol Bagchi, Peeter Kirs, Robert Cerveny, Edward J. Ferrara, Frederick G. Volpicelli, Diomidis Spinellis, and C. Augusto Casas. Forum: Stress fundamentals in CS education, simplicity in production; no room for bias when covering the copyright system; don’t misuse or misteach UML; cover to cover; future CS course already here; to attract new members, increase ACM scope and mandate. Communications of the ACM, 49(8):11– 13, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spinellis:2006:BB

[77] Diomidis Spinellis. Bug busters. IEEE Software, 23(2):92–93, March/April 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:CPL

[78] Diomidis Spinellis. Choosing a programming language. IEEE Software, 23 (4):62–63, July/August 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:DLF

[79] Diomidis Spinellis. Debuggers and logging frameworks. IEEE Software, 23 (3):98–99, May/June 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:OSP

[80] Diomidis Spinellis. Open source and professional advancement. IEEE Software, 23(5):70–71, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:PAP

[81] Diomidis Spinellis. Project asset portability. IEEE Software, 23(1):100– 101, January/February 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2006:RSS

[82] D. Spinellis. Review of ‘Sustainable Software Development: An Agile Perspective’ by Kevin Tate, Addison-Wesley Professional, 2005, $39.99, ISBN 0-321-28608-1. ACM Queue: Tomorrow’s Computing Today, 4(10):49, December/ January 2006–2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [125].

Spinellis:2006:STM

[83] Diomedis Spinellis. Some types of memory are more equal than others. ;login: the USENIX Association newsletter, 31(2):29–36, April 2006. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.usenix.org/
                                        publications/login/2006-04/pdfs/
                                        spinellis.pdf.

Spinellis:2007:AV

[84] Diomidis Spinellis. Abstraction and variation. IEEE Software, 24(5):24– 25, September/October 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:B

[85] Diomidis Spinellis, Naseem Mariam, and Caroline Pepa. Bookshelf. IEEE Software, 24(4):107–109, July/August 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

                                        tronic). URL http://csdl.computer.

org/comp/mags/so/2007/04/s4107.
pdf.

Spinellis:2007:CSR

[86] Diomidis Spinellis. Cracking software reuse. IEEE Software, 24(1):12–13, January/February 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:FSV

[87] Diomidis Spinellis and Panagiotis Louridas. A framework for the static verification of api. The Journal of Systems and Software, 80(7):1156–1168, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Spinellis:2007:P

[88] Diomidis Spinellis. On paper. IEEE Software, 24(6):24–25, November/ December 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:S

[89] Diomidis Spinellis. I spy. IEEE Software, 24(2):16–17, March/April 2007. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:SBO

[90] Diomidis Spinellis. Silver bullets and other mysteries. IEEE Software, 24 (3):22–23, May/June 2007. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2007:TWU

[91] Diomidis Spinellis. The tools we use. IEEE Software, 24(4):20–21, July/August 2007. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic).

Helsen:2008:WMJ

[92] Simon Helsen, Arthur Ryman, and Diomidis Spinellis. Where’s my jetpack? IEEE Software, 25(5):18–21, September/October 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Karakoidas:2008:FJO

[93] Vassilios Karakoidas and Diomidis Spinellis. FIRE/J — optimizing regular expression searches with generative programming. Software—Practice and Experience, 38(6):557–573, May ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Louridas:2008:PLS

[94] Panagiotis Louridas, Diomidis Spinellis, and Vasileios Vlachos. Power laws in software. ACM Transactions on Software Engineering and Methodology, 18 (1):2:1–2:??, September 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Spinellis:2008:AMC

[95] Diomidis Spinellis. The Antikythera mechanism: A computer science perspective. Computer, 41(5):22–27, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spinellis:2008:COK

[96] Diomidis Spinellis and Panagiotis Louridas. The collaborative organization of knowledge. Communications of the ACM, 51(8):68–73, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spinellis:2008:RM

[97] Diomidis Spinellis. Rational metaprogramming. IEEE Software, 25(1):78– 79, January/February 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2008:SB

[98] Diomidis Spinellis. Software builders. IEEE Software, 25(3):22–23, May/June 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2008:UAX

[99] Diomidis Spinellis. Using and abusing XML. IEEE Software, 25(2):88–89, March/April 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2008:WWP

[100] Diomidis Spinellis. The way we program. IEEE Software, 25(4):89–91, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Papadopoulos:2009:ADD

[101] Chrissoleon T. Papadopoulos, Michael E. J. O’Kelly, Michael J. Vidalis, and Diomidis Spinellis. Analysis and design of discrete part production lines, volume 31 of Springer Optimization and Its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-89493-4. xx + 279 pp.

Spinellis:2009:BA

[102] Diomidis Spinellis and Georgios Gousios, editors. Beautiful architecture. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15578-6. xix + 404 pp. LCCN QA76.754 .B43 2009. URL http:

//proquest.safaribooksonline.com/
9780596155780.

Spinellis:2009:BET

[103] Diomidis Spinellis. Basic etiquette of technical communication. IEEE Software, 26(6):86–87, November/December 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:JS

[104] Diomidis Spinellis. Job security. IEEE Software, 26(5):14–15, September/October 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:TTD

[105] Diomidis Spinellis. Tools of the trade: Drawing tools. IEEE Software, 26 (3):12–13, May/June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:TTS

[106] Diomidis Spinellis. Tools of the trade: Start with the most difficult part. IEEE Software, 26(2):70–71, March/April 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2009:ULO

[107] Diomidis Spinellis. User-level operating system transactions. Software— Practice and Experience, 39(14):1215– 1233, September 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sprinkle:2009:FDS

[108] Jonathan Sprinkle, Marjan Mernik, Juha-Pekka Tolvanen, and Diomidis Spinellis. FOCUS: Domain-specific modeling: Guest Editors’ introduction: What kinds of nails need a domainspecific hammer? IEEE Software, 26 (4):15–18, July/August 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Androutsellis-Theotokis:2010:MBA

[109] Stephanos Androutsellis-Theotokis, Diomidis Spinellis, Panos Louridas, and Kostas Stroggylos. A market-based approach to managing the risk of peer-to-peer transactions. Computer Networks (Amsterdam, Netherlands: 1999), 54(5): 675–688, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Spinellis:2010:CD

[110] Diomidis Spinellis. Code documentation. IEEE Software, 27(4):18–19, July/August 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2010:CRB

[111] Diomidis Spinellis. CScout: A refactoring browser for C. Science of Computer Programming, 75(4):216–231, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Spinellis:2010:FD

[112] Diomidis Spinellis. Farewell to disks.

IEEE Software, 27(6):82–83, November/ December 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2010:ST

[113] Diomidis Spinellis. Software tracks. IEEE Software, 27(2):10–11, March/ April 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2010:UE

[114] Diomidis Spinellis. UML everywhere. IEEE Software, 27(5):90–91, September/October 2010. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:AD

[115] Diomidis Spinellis. Agility drivers. IEEE Software, 28(4):96, 95, July/August 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:CUO

[116] Diomidis Spinellis. Choosing and using open source components. IEEE Software, 28(3):96, 95, May/June 2011. CO-DEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:EE

[117] Diomidis Spinellis. elytS edoC. IEEE Software, 28(2):104, 103, March/April 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2011:FI

[118] Diomidis Spinellis. Faking it. IEEE Software, 28(5):96, 95, September/October 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gutknecht:1994:PLS

[119] Jürg Gutknecht, editor. Programming languages and system architectures: international conference, Zurich, Switzerland, March 2–4, 1994: proceedings, volume 782 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 0-387-57840-4 (New York), 3-540-57840-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .P79 1994. DM72.00.

Karagiannis:1994:DES

[120] D. Karagiannis, editor. Database and expert systems applications: 5th international conference, DEXA ’94, Athens, Greece, September 7–9, 1994: proceedings, volume 856 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 0-387-58435-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D35843 1994.

USENIX:1994:PWU

[121] USENIX Association, editor. Proceedings of the Winter 1994 USENIX Conference: January 17–21, 1994, San Francisco, California, USA. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

USENIX:1997:PCD

[122] USENIX, editor. Proceedings of the Conference on Domain-Specific Languages, October 15–17, 1997, Santa Barbara, California. USENIX Association, Berkeley, CA, USA, 1997. ISBN 1-880446-89-8. LCCN QA76.7 .C663

1997. URL http://www.usenix.org/
publications/library/proceedings/
dsl97/.

USENIX:2000:UAT

[123] USENIX, editor. 2000 USENIX Annual Technical Conference: San Diego, CA, USA, June 18–23, 2000. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-22-7. LCCN

                                        ???? URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix2000.

USENIX:2001:PFT

[124] USENIX, editor. Proceedings of the FREENIX Track: 2001 USENIX Annual Technical Conference: June 25–30, 2001, Marriott Copley Place Hotel, Boston, Massachusetts, USA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL http://

                                        www.usenix.org/publications/library/
                                        proceedings/usenix01/freenix01/
                                        technical.html.

Tate:2006:SSD

[125] Kevin Tate. Sustainable software development: an agile perspective. Agile software development series. Addison-Wes-ley, Reading, MA, USA, 2006. ISBN 0-321-28608-1. xxxiii + 226 pp. LCCN QA76.76.D47 T38 2006. URL http:/

                                        /www.loc.gov/catdir/toc/ecip0515/
                                        2005019120.html.